4. The impulse response of a system is given by h[n]=(0.3)u[n]. If the input to the system is x[r]=(-0.6) u [n], giving an
Answer #1

\mathbf{Given\:,\:}

\mathbf{\:h[n]\:=\:(0.3)^n\,u[n]\:\:\:and\:\:\:x[n]\:\:=\:\:(-0.6)^n\,u[n]}

\mathbf{y[n]\:\:=\:\:x[n]\ast h[n]\:\:\:\:}

\mathbf{\Rightarrow\:\:\:Y(e^{j2\pi f})\:=\:X(e^{j2\pi f}) \cdot H(e^{j2\pi f})}

\mathbf{(a)}

\mathbf{\underline{NOTE:}\: \: \: \: \: \: \: a^n\,u[n]\: \: \overset{F.T}{\longleftrightarrow}\:\:\frac{1}{1-a\,e^{-j2\pi f}}}

\mathbf{Then\: ,\:\: \: \: \: H(e^{j2\pi f})\:=\frac{1}{1-0.3\,e^{-j2 \pi f}}}

\mathbf{X(e^{j2\pi f})\:=\frac{1}{1+0.6\,e^{-j2 \pi f}}}

\mathbf{Therefore\: ,\:\: Y(e^{j2\pi f})\:=\left ( \frac{1}{1-0.3\,e^{-j2 \pi f}} \right )\left ( \frac{1}{1+0.6\,e^{-j2 \pi f}} \right )}

\mathbf{(b)}

\mathbf{Y(e^{j2\pi f})\:\:\:=\: \: \left ( \frac{1}{1-0.3\,e^{-j2 \pi f}} \right )\left ( \frac{1}{1+0.6\,e^{-j2 \pi f}} \right )}

\mathbf{=\: \:\frac{1}{3}\left ( \frac{1}{1-0.3\,e^{-j2 \pi f}}\: \: +\: \: \frac{2}{1+0.6\,e^{-j2 \pi f}} \right )}

\mathbf{Now\: \: by \: \: applying \: \: Inverse\: \: DTFT\: \: to\: \: Y(e^{j2\pi f})\:,}

\mathbf{y[n]\:\:=\:\:\frac{1}{3}\left [ \:0.3^n+2(-0.6)^n \right ]u[n]}